Given that sinx is a solution of the differential equation: L[y]=d^4y/dx^4+2d^3y/dx^3+6d^2ydx^2+2dy/dx+5y=0 and the 4 parameter family of solutions of it. Hence solve L[y]=x^2
1
Expert's answer
2020-08-11T18:28:55-0400

"Solution"

To evaluate the particular solution "l[y]=x^2" .

It was given that "\sin(x)" is a solution of


"\frac{d^4y}{dx^4}+2\frac{d^3y}{dx^3}+6\frac{d^2y}{dx^2}+2\frac{dy}{dx}+5y=0"

Next,


"\frac{d^4y}{dx^4}+2\frac{d^3y}{dx^3}+6\frac{d^2y}{dx^2}+2\frac{dy}{dx}+5y=x^2"

For D.E "\frac{d^4y}{dx^4}+2\frac{d^3y}{dx^3}+6\frac{d^2y}{dx^2}+2\frac{dy}{dx}+5y=x^2" , you need to try the solution

"y_{p_1}=Ax^2+Bx+C" (Since "l[y]=x^2" is of degree 2), Such that


"\frac{d^2y}{dx^2}=A, \frac{dy}{dx}=B"


"\frac{d^3y}{dx^3}=\frac{d^4y}{dx^4}=0"

"6A+2B+5(Ax^2+Bx+c)=x^2"

"5Ax^2+6A+5Bx+2B+5C=x^2"


Equating the coefficients of like powers yields:



"5A=1 \implies A=\frac{1}{5}"

"5B=0 \implies B=0"

"6A+2B+5C=0 \implies 6A+5C=0 \implies 5C=-\frac{6}{5} \implies C=-\frac{6}{25}"

Hence, evaluating "y_{p_1}" yields


"y_{p_1}=\frac{1}{5}x^2-\frac{6}{25}"